Stephanie's Bridal Shoppe sells wedding dresses. The average selling price of each dress is $1,000, variable costs are $400, and fixed costs are $90,000. How many dresses must the Bridal Shoppe sell to yield after-tax net income of $18,000, assuming the tax rate is 40%? A) 200 dresses B) 170 dresses C) 150 dresses D) 145 dresses
Answer #1
Required before tax income = 18000/(1-40%)= $30000
Number of dresses to be sold = (Fixed costs+Before tax income)/Unit Contribution margin
Number of dresses to be sold = (90000+30000)/(1000-400)= 200 dresses
Option A is correct